# Enhancing LLM-Based Short Answer Grading with Retrieval-Augmented Generation

### v1

@@ 1 Introduction
Assessment and analysis of student understanding in science education extend far beyond simple grading. With the emergence of new frameworks for K-12 science education, there is an increasing emphasis on analyzing students’ multidimensional comprehension of scientific understanding. The National Research Council’s Framework for K-12 Science Education Council et al. (2012) established three critical dimensions of science learning: disciplinary core ideas (DCIs), science and engineering practices (SEPs), and crosscutting concepts (CCCs). To demonstrate their achievement, students should apply their 3D understanding to interpret compelling phenomena or solve real-world problems. This framework has fundamentally transformed how we assess student learning, shifting away from traditional multiple-choice questions toward short-answer assessments that better capture students’ authentic understanding He et al. (2023b). Short-answer questions have become particularly valuable in this context, as they require students to demonstrate their three-dimensional knowledge by explaining real-world phenomena and solving complex problems He et al. (2023a). These responses provide rich insights into students’ conceptual understanding, revealing both their mastery and misunderstanding, which is essential for teachers to adapt their instructional strategies and support students’ self-regulated learning He et al. (2024).
@@
However, the analysis of these short answers presents significant challenges as it is time-intensive, requires deep expertise across all three dimensions, and makes it difficult for teachers to provide timely, actionable feedback to individual students. Fortunately, artificial intelligence (AI), particularly automatic short answer grading (ASAG) systems, has emerged as a promising solution to these challenges. Traditional ASAG approaches using machine learning (ML) techniques have shown success in providing consistent and objective scoring but are limited by their reliance on large training datasets Burrows et al. (2015); Sultan et al. (2016). The recent advent of large language models (LLMs) has opened new possibilities for training ASAG systems on smaller datasets, due to its possession of broad general knowledge from pre-training. However, current LLM-based approaches face two critical limitations in science education assessment. First, they exhibit deceptive performance in technical domains, often generating plausible-sounding but scientifically false evaluations when in-depth domain knowledge is required. Their output can be unreliable when evaluating answers that contain ambiguous terminology or complex scientific concepts. Second, they lack sufficient task-specific knowledge about the grading criteria Lewis et al. (2020). They often fail to understand the nuanced requirements of specific educational frameworks such as the three-dimensional learning approach, which can lead to misalignment between automated scores and expert evaluations.
@@
To address these limitations, we introduce GradeRAG, a novel framework that enhances LLM-based grading through retrieval-augmented generation (RAG) Zeng et al. (2024). Our approach implements a specialized RAG pipeline that provides the LLM with access to curated domain-specific knowledge bases, enabling a more accurate assessment of scientific concepts across the three dimensions (DCIs, SEPs, and CCCs). This knowledge-focused approach is particularly crucial in science education, where effective assessment requires domain-specific expertise of specialized concepts and their interconnections He et al. (2023a); Beatty et al. (2014). As part of our knowledge retrieval strategy, we incorporate expert-annotated scoring rationales as a specialized form of knowledge source. These examples contain detailed scoring rationales that explicitly identify critical scientific terminology and reasoning patterns in student responses. By treating these examples as retrievable knowledge, LLMs are guided to emulate expert analysis processes, further improving scoring performance. We evaluate GradeRAG on a dataset of student short answers on science assessments. Results demonstrate improvements in both grading accuracy and consistency compared to the baseline approaches. Our results suggest that integrating specialized knowledge retrieval systems can bridge the gap between automated efficiency and expert-level assessment in science education, leading to more reliable evaluation of complex scientific understanding.
@@ 2 Related Work
Automatic short answer grading (ASAG) has evolved significantly over the past decades. Early approaches mainly focus on text matching and statistics-based methods Mohler et al. (2011). With the advance in ML, later systems employ feature engineering techniques with supervised training methods Leacock and Chodorow (2003). These traditional approaches typically require large training datasets and often struggle with capturing the semantic differences in student answers. Recent works utilizing neural networks demonstrate improvement in understanding semantic relationships in short answers Hassan et al. (2018). Large language models (LLMs) further enable enhanced capabilities in ASAG through zero-shot and few-shot Lee et al. (2024) techniques. However, as noted by Li et al. (2023), LLMs still face challenges in specialized domains like science education, where domain-specific knowledge is crucial for accurate assessment.
@@
Retrieval-augmented generation (RAG) has emerged as an effective approach to enhance LLMs’ reasoning ability by providing access to external knowledge sources. RAG has demonstrated effectiveness in knowledge-intensive tasks requiring factual accuracy Lewis et al. (2020) and domain expertise Guu et al. (2020). In the educational context, the application of RAG remains relatively unexplored, with most works focusing on content generation Miladi et al. (2024) rather than assessment. Recent studies have begun investigating RAG for automated assessment. For example,  Harshavardhan and Singh (2024) proposes a rubric-centric approach for automated test correction using RAG, while  Sundar et al. (2024) evaluated RAG frameworks for grading open-ended written responses. However, the application of RAG, specifically for short answer grading in science education, where domain-specific knowledge across multiple dimensions (DCIs, SEPs, CCCs) is crucial, represents a novel contribution to our work.
@@ 3.2 RAG-Based Grading System
Illustration of the proposed framework Refer to caption
@@ 3.2.1 Knowledge Base Construction
As shown in Figure  1, our system constructs a triple-index knowledge base optimized for scientific assessment. $K_{D_{1}}$ and $K_{D_{2}}$ consists of project-related information and grading guidelines, while $K_{E}$ (which is optinal) contains expert-annotated scoring examples. The first index $K_{D_{1}}$ contains general assessment documents such as the Framework of K-12 Science Education Council et al. (2012) and the NGSS standard document States (2013). These materials are systematically chunked into coherent segments of approximately 512 tokens, which balances content completeness with the retrieval granularity. This chunking size preserves sufficient context while enabling precise retrieval of relevant concepts, as smaller chunks would fragment scientific explanations while larger chunks can introduce excessive irrelevant information. Each chunk is tagged with metadata including the document source and dimension type to facilitate efficient filtering during retrieval.
The second index $K_{D_{2}}$ consists of task-specific materials, such as the three-dimensional learning progression materials He et al. (2023a), which includes unpacking DCIs, SEPs, and CCCs materials and the holistic levels in the three assessment tasks. For these more specialized resources, we employ a finer-grained, manual chunking strategy. Specifically, we divide the content into single sentences or meaningful segments containing one or two bullet points. This approach is adopted as task-specific materials contain dense, concentrated information where even small segments can provide critical assessment criteria. With this fine-grained chunking strategy, our system can retrieve precisely the relevant assessment standards without including extraneous information. Each chunk is tagged with metadata indicating the dimension type, level number, and chunk identifier to enable targeted filtering during retrieval.
@@ 3.2.3 Grading
Figure 3: An example prompt for grader

### v2

@@ 1 Introduction
Assessment and analysis of student understanding in science education extend far beyond simple grading. With the emergence of new frameworks for K-12 science education, there is an increasing emphasis on analyzing students’ multidimensional comprehension of scientific understanding. The National Research Council’s Framework for K-12 Science Education national2012framework established three critical dimensions of science learning: disciplinary core ideas (DCIs), science and engineering practices (SEPs), and crosscutting concepts (CCCs). To demonstrate their achievement, students should apply their 3D understanding to interpret compelling phenomena or solve real-world problems. This framework has fundamentally transformed how we assess student learning, shifting away from traditional multiple-choice questions toward short-answer assessments that better capture students’ authentic understanding he2023predicting . Short-answer questions have become particularly valuable in this context, as they require students to demonstrate their three-dimensional knowledge by explaining real-world phenomena and solving complex problems he2023applying . These responses provide rich insights into students’ conceptual understanding, revealing both their mastery and misunderstanding, which is essential for teachers to adapt their instructional strategies and support students’ self-regulated learning he2024design .
@@
However, the analysis of these short answers presents significant challenges as it is time-intensive, requires deep expertise across all three dimensions, and makes it difficult for teachers to provide timely, actionable feedback to individual students. Fortunately, artificial intelligence (AI), particularly automatic short answer grading (ASAG) systems, has emerged as a promising solution to these challenges. Traditional ASAG approaches using machine learning (ML) techniques have shown success in providing consistent and objective scoring but are limited by their reliance on large training datasets burrows2015eras ; sultan2016fast . The recent advent of large language models (LLMs) has opened new possibilities for training ASAG systems on smaller datasets, due to its possession of broad general knowledge from pre-training. However, current LLM-based approaches face two critical limitations in science education assessment. First, they exhibit deceptive performance in technical domains, often generating plausible-sounding but scientifically false evaluations when in-depth domain knowledge is required. Their output can be unreliable when evaluating answers that contain ambiguous terminology or complex scientific concepts. Second, they lack sufficient task-specific knowledge about the grading criteria lewis2020retrieval . They often fail to understand the nuanced requirements of specific educational frameworks such as the three-dimensional learning approach, which can lead to misalignment between automated scores and expert evaluations.
@@
To address these limitations, we introduce GradeRAG, a novel framework that enhances LLM-based grading through retrieval-augmented generation (RAG). Our approach implements a specialized RAG pipeline that provides the LLM with access to curated domain-specific knowledge bases, enabling a more accurate assessment of scientific concepts across the three dimensions (DCIs, SEPs, and CCCs). This knowledge-focused approach is particularly crucial in science education, where effective assessment requires domain-specific expertise of specialized concepts and their interconnections he2023applying ; beatty2014developing . As part of our knowledge retrieval strategy, we incorporate expert-annotated scoring rationales as a specialized form of knowledge source. These examples contain detailed scoring rationales that explicitly identify critical scientific terminology and reasoning patterns in student responses. By treating these examples as retrievable knowledge, LLMs are guided to emulate expert analysis processes, further improving scoring performance. We evaluate GradeRAG on a dataset of student short answers on science assessments. Results demonstrate improvements in both grading accuracy and consistency compared to the baseline approaches. Our results suggest that integrating specialized knowledge retrieval systems can bridge the gap between automated efficiency and expert-level assessment in science education, leading to more reliable evaluation of complex scientific understanding.
@@ 2 Related Work
Automatic short answer grading (ASAG) has evolved significantly over the past decades. Early approaches mainly focus on text matching and statistics-based methods mohler2011learning . With the advance in ML, later systems employ feature engineering techniques with supervised training methods leacock2003c-rater . These traditional approaches typically require large training datasets and often struggle with capturing the semantic differences in student answers. Recent works utilizing neural networks demonstrate improvement in understanding semantic relationships in short answers hassan2018automatic . Large language models (LLMs) further enable enhanced capabilities in ASAG through zero-shot and few-shot lee2024applyingllm techniques. However, as noted by li2023adapting , LLMs still face challenges in specialized domains like science education, where domain-specific knowledge is crucial for accurate assessment.
@@
Retrieval-augmented generation (RAG) has emerged as an effective approach to enhance LLMs’ reasoning ability by providing access to external knowledge sources. RAG has demonstrated effectiveness in knowledge-intensive tasks requiring factual accuracy lewis2020retrieval and domain expertise guu2020retrieval . In the educational context, the application of RAG remains relatively unexplored, with most works focusing on content generation miladi2024comparative rather than assessment. Recent studies have begun investigating RAG for automated assessment. For example,  harshavardhan2024rubric proposes a rubric-centric approach for automated test correction using RAG, while  sundar2024revolutionizing evaluated RAG frameworks for grading open-ended written responses. However, the application of RAG, specifically for short answer grading in science education, where domain-specific knowledge across multiple dimensions (DCIs, SEPs, CCCs) is crucial, represents a novel contribution to our work.
@@ 3.2 RAG-Based Grading System
Illustration of the proposed framework
@@ 3.2.1 Knowledge Base Construction
As shown in Figure  1, our system constructs a triple-index knowledge base optimized for scientific assessment. $K_{D_{1}}$ and $K_{D_{2}}$ consists of project-related information and grading guidelines, while $K_{E}$ (which is optinal) contains expert-annotated scoring examples. The first index $K_{D_{1}}$ contains general assessment documents such as the Framework of K-12 Science Education national2012framework and the NGSS standard document ngss2013next . These materials are systematically chunked into coherent segments of approximately 512 tokens, which balances content completeness with the retrieval granularity. This chunking size preserves sufficient context while enabling precise retrieval of relevant concepts, as smaller chunks would fragment scientific explanations while larger chunks can introduce excessive irrelevant information. Each chunk is tagged with metadata including the document source and dimension type to facilitate efficient filtering during retrieval.
The second index $K_{D_{2}}$ consists of task-specific materials, such as the three-dimensional learning progression materials he2023applying , which includes unpacking DCIs, SEPs, and CCCs materials and the holistic levels in the three assessment tasks. For these more specialized resources, we employ a finer-grained, manual chunking strategy. Specifically, we divide the content into single sentences or meaningful segments containing one or two bullet points. This approach is adopted as task-specific materials contain dense, concentrated information where even small segments can provide critical assessment criteria. With this fine-grained chunking strategy, our system can retrieve precisely the relevant assessment standards without including extraneous information.
Each chunk is tagged with metadata indicating the dimension type, level number, and chunk identifier to enable targeted filtering during retrieval.
@@ 3.2.3 Grading
Figure 3: An exemplary prompt for grader
@@ 5 Conclusion
6 Acknowledgments
This study is supported by the National Science Foundation [grant number DRL-2446701]. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.